### v2

# Penetrance Estimation in Family-based Studies with the penetrance R package

Abstract
Reliable methods for penetrance (age-specific risk among those who carry a genetic variant) estimation are critical to improving clinical decision making and risk assessment for hereditary syndromes. We introduce penetrance, an open-source R package, to estimate age-specific penetrance using family-history pedigree data. The package employs a Bayesian estimation approach, allowing for the incorporation of prior knowledge through the specification of priors for the parameters of the carrier distribution. It also includes options to impute missing ages during the estimation process, addressing incomplete age information which is not uncommon in pedigree datasets. Our software provides a flexible and user-friendly tool for researchers to estimate penetrance in complex family-based studies, facilitating improved genetic risk assessment in hereditary syndromes. The penetrance R package is freely available on CRAN, with accompanying documentation and examples.
1 Introduction
Precision prevention and early detection of heritable diseases rely on identifying individuals at increased risk. Comprehensive germline DNA analysis panels can reliably detect pathogenic germline variants (PGVs) that may increase disease risk. In this context, an accurate estimate of age-specific risk of disease is a crucial input for counseling individuals.
In this paper, a “PGV” refers specifically to a germline allele known or strongly suspected to be pathogenic. Risk counseling can be performed at the level of the specific PGV, but more often the interest lies in estimating risk for groups of PGVs that occur in the same gene, gene region, or group of genes. For example, one may be interested in the risk associated with the I307K mutation in the APC gene, any PGV in the APC gene, or any PGV in a group of genes comprising Lynch syndrome. Here, we define a ”carrier” as an individual who has a PGV within a specific set of interest. Thus our model can be applied equally to the estimation of PGV-specific, gene-specific or syndrome-specific penetrance.
A penetrance function is the proportion of individuals who develop a condition by a certain age among carriers. Because of the large number of genes probed by genetic panel testing, the relative rarity of PGVs in many of them, and the importance of investigating risk among specific ethnic subgroups, as well as penetrance modifiers, penetrance estimation is emerging as a critical bottleneck in implementing precision prevention for individuals at increased familial risk.
Penetrance estimation has been extensively studied in the past for a variety of study designs (Thomas, 2004). Family-based studies usually center around a proband who is typically the initial point of contact with the healthcare provider and is the individual reporting their family history. The proband is often the only genotyped individual in the family and may or may not report additional genotyping information for known relatives. Different methods for penetrance estimation have also been previously developed for family-based studies (Chen et al., 2009; Shin et al., 2020). Classical segregation analysis, for example, has been widely used to infer inheritance patterns and estimate penetrance, but it often requires large, well-characterized pedigrees and can yield imprecise estimates when sample sizes are small (Elston and Stewart, 1971; Claus et al., 1991).
However, in our assessment, there is a lack of easily usable software for penetrance estimation for family-based studies that implements a Bayesian estimation approach. For example, the widely used software package MENDEL by Lange et al. (2001, 2008, 2016) offers an option for penetrance estimation which can account for a full pedigree, but is difficult to use and interpret. Other penetrance estimation software and webtools are outdated and no longer maintained. To our knowledge no other software packages for penetrance estimation are readily available today. Additionally, when analyzing a study of PGV carriers, it can be important to integrate prior knowledge from published studies, which provide important information. Currently available software packages do not provide options to easily include prior knowledge.
We introduce an approach and the associated open-source R software package, called penetrance, to carry out age-specific and sex-specific penetrance estimation for complex family-based studies where carrier status is not observed for all family members. The implementation is based on Bayesian estimation methods, which allows for the incorporation of prior knowledge from existing studies and provides a flexible and easy-to-use method for penetrance estimation. The penetrance function for carriers is modeled using a modified Weibull distribution.
While this approach has broad applications across many heritable conditions, our primary motivation comes from cancer genetics, where early detection and prevention strategies have shown significant impact on patient outcomes (Rebbeck et al., 2018, 2009; Finch et al., 2006). In the following we introduce the software and provide the user with an illustration of our penetrance estimation approach and package, applying it to estimate colorectal cancer penetrance among MLH1 carriers using a simulation study motivated by real data of 130 probands with PGVs in the MLH1 gene, reporting family history on 4,604 individuals (including probands).
2 Methods
2.1 Model
We consider a pedigree-based framework for analyzing disease risk. While some sets of PGVs may increase risk for multiple diseases, our framework focuses exclusively on analyzing the age-specific penetrance for a single disease at a time. In our model, families are indexed by $i=1,\dots,I$, with members indexed by $j=0,\dots,J_{i}$ ($j=0$ denotes the proband). For family $i$, disease status are represented by $H_{i}=(H_{i0},\dots,H_{iJ_{i}})$, where each $H_{ij}=(d_{ij},y_{ij})$ denotes individual $j$’s binary diagnosis status ($d_{ij}$) and age at diagnosis or censoring ($y_{ij}$) for a single disease.
An individual’s carrier status is denoted by $G_{ij}$, which takes a binary value to indicate whether individual $i$ is a carrier ($G_{ij}=1$) or not ($G_{ij}=0$). We assume here that homozygous carriers of the PGV are not viable. The model accommodates missing carrier status data, with $G_{i,\text{obs}}$ and $G_{i,\text{mis}}$ denoting observed and missing carrier status respectively, such that $G_{i}=(G_{i,\text{obs}},G_{i,\text{mis}})$. We derive the family-level likelihood:
$P(H_{i}\mid G_{i,\text{obs}})=\sum_{G_{i,\text{mis}}}P(H_{i}\mid G_{i})\,P(G_{i,\text{mis}}\mid G_{i,\text{obs}}).$
Our methodology is grounded in fundamental genetic and statistical assumptions. The genetic model relies on critical population-level parameters, including the carrier prevalence in the general population and the baseline age-specific probabilities of developing the disease for noncarriers. The prevalence would be approximately two times the allele frequency.
We assume conditional independence of individual disease status given the complete set of carrier status for the pedigree, and assume equivalent disease status among distinct pathogenic variants within the same gene. The rare disease syndromes are modeled with dominant inheritance patterns and focus on the first occurrence of the disease in an individual. In particular, cancer syndromes mostly exhibit dominant inheritance patterns (Shin et al., 2020; King et al., 2003; Lynch and de la Chapelle, 1999). These assumptions enable the factorization of the likelihood into individual-specific components:
$P(H_{i}\mid G_{i})=\prod_{j=0}^{J_{i}}P(H_{ij}\mid g_{ij}).$
We implement the likelihood computation using the Elston-Stewart peeling algorithm through the clipp package, facilitating efficient calculation of carrier probabilities in complex pedigree structures (Golicher, 2023; Elston and Stewart, 1971).
2.2 Weibull Parameterization
We choose a Weibull distribution for our parametric estimation approach, extending it with threshold and asymptote parameters to account for minimum onset age and lifetime disease probability. The standard Weibull distribution with scale ($\alpha>0$) and shape ($\beta>0$) parameters is modified by introducing a threshold parameter ($\delta>0$), which shifts the distribution along the age axis, and an asymptote parameter ($10<\gamma<$ 1), representing the lifetime probability of developing the disease. The model’s parameters $\theta=(\alpha,\beta,\gamma,\delta)$ define the cumulative distribution function (CDF) that represents disease onset probability over time. The model can incorporate prior information through informative prior distributions, which can be specified using more interpretable conditional quartiles (see Appendix A).
2.3 Bayesian Estimation Approach
We employ a parametric Bayesian approach to find the posterior distribution of the parameters of the specified Weibull distribution, $\theta=(\alpha,\beta,\gamma,\delta)$. The direct calculation of this posterior is intractable due to the complexity of the likelihood function, which involves modeling dependencies within families and integrating over unobserved variables. Therefore, we utilize MCMC methods to approximate the posterior distribution. MCMC methods generate samples from a target distribution by constructing a Markov chain whose stationary distribution is the desired posterior distribution $P(\theta\mid H)$ (Metropolis et al., 1953; Gilks et al., 1995). Our approach uses a fine-tuned adaptive MCMC algorithm that ensures adequate exploration of the parameter space.
3 Package
The methods discussed above are implemented in an easy-to-use R software package, available on CRAN. The package workflow includes three main parts: (1) the user input, including family data in the form of pedigrees, and the specifications of the options for the penetrance estimation, (2) the estimation of the posterior distribution using the MCMC approach, and (3) the output of the results in the form of the samples from the approximated posterior distribution, i.e. the estimated penetrance function. Figure 1 shows the workflow.
Figure 1: Workflow of the penetrance Package with dependency on clipp.
Table 1: List of main functions in penetrance.
Name Category Description
makePriors Pre-processing Creates prior distributions based on the user inputs. Depending on the combination of user inputs, the default setting for the prior distributions are modified.
mhChain Algorithm Estimates the penetrance probability using the MCMC approach. The inputs include the formatted family pedigree data, options for the construction of the MCMC chains, and the output of makePriors.
mhLoglikelihood Algorithm Evaluates the the log-likelihood for the given data and model parameters using the clipp package for the likelihood calculation.
penetrance Main Function Runs the main function. The inputs include the formatted family pedigree data, options for the construction of the MCMC chains, and some options for the output. The default output is a list of the posterior samples for each of the parameters from each of the independently run chains.
3.1 User Input
3.1.1 Pedigree Input
The main input is the family pedigree data reported by the proband. The pedigree file is an R data frame object. Each family tree is represented by a single data frame that includes the following columns: PedigreeID, ID, Sex, MotherID, FatherID, isProband, CurAge, isAff, Age, and Geno. Each row in the data frame corresponds to a unique family member, identified by a unique ID within the family and a unique PedigreeID. Connections between family members are established through the MotherID and FatherID columns, which reference the IDs of each individual’s mother and father, if this information is available. In each family tree data frame, one individual is designated as the proband indicated by the isProband column with a value of 1. In the type of family-based penetrance studies considered here, families are typically ascertained through probands who are confirmed carriers. The column CurAge indicates the censoring age, which is the current age if the person is alive or the age at death if the person is deceased, with allowed ages ranging from 1 to $max\_age$. Missing ages are allowed and can be coded as NA or left empty. The isAff column indicates the disease affection status, with 1 for diagnosed individuals, 0 for unaffected, and NA for unknown. The Age column indicates the age of disease diagnosis, encoded as NA if the individual was not diagnosed or if the age was unknown, with allowed ages ranging from 1 to $max\_age$. The Geno column includes germline testing results, i.e. the carrier status, with positive results coded 1, negative results coded 0, and unknown results as NA or left empty. Every family tree must include a PedigreeID, a numeric value representing the unique identifier for each family, ensuring that there are no duplicate entries. The Sex field must be coded as 0 for females or 1 for males, with unknown sex coded as NA. More details on the structure of the pedigree required for penetrance is available in the Fam3Pro (formerly PanelPRO) documentation where a similar structure is used (Lee et al., 2021).
3.1.2 Prior Specification
To run the MCMC algorithm described in the methods section, the prior distributions for $\theta=(\alpha,\beta,\gamma,\delta)$ need to be specified. The package provides the user with a flexible approach to prior specification, balancing customization with an easy-to-use workflow. In addition to providing default prior distributions (setting 1), the package allows users to customize the priors directly (setting 2) or based on penetrance estimates from the literature (setting 3). The following settings for the prior distribution specification are available:
  1. 1.
    Default: The default setting employs pre-specified, uninformative priors. The default setting for the priors can be seen in Table 2.
  2. 2.
    Custom Parameters: The user has the option to tune the parameters of the default setting directly by overwriting the parameters in the $prior\_params\_default$ object (see Figure 2).
  3. 3.
    Prior Elicitation from Existing Studies: The package provides additional functionality to facilitate in the prior elicitation based on existing penetrance studies. For more details on the way options to input information from published studies, see Appendix A.
(a) Default prior parameters (Setting 1)
(b) Example configuration with user-modified prior parameters (Setting 2)
Figure 2: Prior Specification.
Model Parameter Default Prior Distribution
First Quartile $Q_{25}$ Scaled Beta(6,3)
Median $Q_{50}$ Scaled Beta(2,2)
Asymptote $\gamma$ Scaled Beta(1,1)
Threshold $\delta$ Uniform(5,30)
Table 2: Default prior distributions implemented in penetrance.
Based on the user inputs, penetrance utilizes the $makePriors$ function to generate priors for each of the four Weibull distribution parameters for both males and females. The function applies the same prior distributions to both sexes, using four unique priors for all eight parameters.
3.1.3 Estimation Model Inputs
The estimation procedure follows the approach outlined in the methods section. In addition to the pedigree data ($data$) and the specified priors ($prior\_params$) the user inputs a carrier prevalence ($prev$) and a baseline age-specific probability of developing the disease ($baseline\_data$). Assuming carriers are rare, the population age-specific probability of developing the disease is a good approximation for the noncarrier penetrance. Such population-level risk data usually comes from a population registry such as the Surveillance, Epidemiology, and End Results (SEER) Program for cancer in the United States (Howlader et al., 2020). Otherwise, the user may choose to calculate the noncarrier penetrance using the calculateNCPen function by setting BaselineNC = TRUE. This enables the calculation of a noncarrier penetrance as part of the estimation process, using the Weibull parameters from the current iteration. Per default, the estimation is sex-specific, estimating two distinct set of parameters for females and males. However, the user has the option to run the estimation without considering sex, by setting sex_specific = FALSE. The user can provide additional inputs to the function at run-time. An overview of the options the user can specify in the main function call can be found in Table 3.
Table 3: List of model options for penetrance (excluding prior and output options).
Argument Name Value Default Value Definition
data NA NA Family pedigree data frame in the appropriate format.
twins List NA Identical twins or triplets in the family can be specified.
n_chains Integer $>$ 0 1 Number of chains for parallel computation.
n_iter_per_chain Integer $>$ 0 10000 Number of iterations per chain.
ncores Integer $>$ 0 6 Number of cores for parallel computation.
baseline_data Dataset Default Object Data for baseline risk estimates.
max_age Integer $>$ 0 94 Maximum age considered for analysis.
remove_proband Boolean FALSE Logical indicating whether to remove probands from the analysis.
age_imputation Boolean FALSE Logical indicating whether to perform age imputation.
imp_interval Integer $>$ 0 10 The interval at which age imputation for affected and unaffected missing ages should be performed when age_imputation = TRUE.
sex_specific Boolean TRUE Logical indicating whether to perform sex-specific estimation.
median_max Boolean TRUE Whether to use baseline median age or max_age as upper bound for median proposal.
BaselineNC Boolean TRUE If TRUE, noncarrier penetrance is assumed to be the baseline penetrance.
var Vector c(0.1, 0.1, 2, 2, 5, 5, 5, 5) Vector of initial variances for the covariance matrix of the proposal distribution.
burn_in Fraction (0 to 1) 0 Fraction of results to discard as burn-in.
thinning_factor Integer $\geq$ 1 1 Thinning factor, applied to individual Markov chains.
distribution_data List or NA Default Object Data used to generate prior distributions.
prev Float 0.0001 Prevalence (of the carrier status) in the population.
3.2 Estimation Algorithm
3.2.1 Preprocessing of Pedigree Data
The estimation requires preparation of the pedigree data, particularly handling missing age information and adjusting the data format for compatibility with the clipp package. Firstly, checks are made to validate the input data and parameters. The $transformDF$ function takes the input data in the pedigree format and transforms it for the use with penetrance and clipp. If the remove_proband parameter is set to TRUE, it removes the disease status of the probands from the data. This approach serves as a strategy for ascertainment correction, though it comes at the cost of excluding probands’ information on disease status from the sample. The appropriateness of this strategy depends on the study design and the specific mechanisms used to ascertain probands into the cohort.
3.2.2 Age Imputation
By default, the package assumes that the age information on probands and relatives are complete and no age imputation is required. In the case where an individual lacks an age of diagnosis or censoring age, their age-related contribution is omitted from the likelihood calculation. To address datasets with missing age data, the penetrance package provides an option to automatically impute missing ages of diagnosis or censoring ages. The imputation is based on the individual’s affection status, sex, and degree of relationship to the proband, who is typically a carrier. We address missing ages of diagnosis in affected individuals by computing posterior carrier probabilities to impute missing ages by sampling from the estimated penetrance function (for likely carriers) or the provided baseline (for likely noncarriers). For unaffected individuals with missing ages, we impute censoring ages from the empirical age distribution of unaffected individuals in the data and update these imputed ages periodically during MCMC to incorporate uncertainty into parameter estimates. The frequency with which ages are imputed affects the run time and can be specified using the option $imp\_interval$. The default is $imp\_interval=10$, which means that age imputation is performed at every 10th iteration.
3.2.3 Algorithm Initialization
The initial set of parameters, denoted $\theta_{0}$, is drawn from the empirical distribution of the provided data, stratified by sex and disease affection status. Specifically, we calculate initial thresholds, medians, and first quartiles for both sexes based on the ages of affected individuals, ensuring that these values fall within the specified prior bounds. Asymptotes for male and female are randomly drawn between the cumulative lifetime risk of the baseline (given by the input parameter $baseline\_data$) and 1.
3.2.4 Sampling of Proposals
In each iteration of the MCMC algorithm, a new proposal is generated for the parameter vector. After generating the proposal, it is essential to ensure that each parameter falls within its respective bounds. These bounds are set to maintain realistic and meaningful parameter values based on biological plausibility (see Appendix B). If any parameter falls outside these bounds, the proposal is automatically rejected, ensuring that the algorithm only considers valid parameter values for further evaluation.
3.2.5 Likelihood Calculation
Based on the proposed parameters of the Weibull distribution, penet.fn calculates the age-specific penetrance for every individual at each iteration of the algorithm. It assumes a genetic model with a single biallelic locus. Age-specific penetrance probabilities are adjusted based on carrier status and affection status of the individual. As mentioned above, we assume that homozygous carriers are not viable.
In every iteration of the algorithm, the (log-)likelihood for the families in the pedigree file is calculated. The likelihood evaluation for a given family and penetrance matrix is performed by the $pedigree\_loglikelihood$ function from the clipp R package.
3.3 Output
The penetrance function returns a comprehensive set of outputs from the MCMC estimation process. Each chain generates its own posterior samples, and these samples collectively represent draws from the target posterior distribution. After running penetrance, the outputs consist of:
  • The posterior samples for the first quartile, median, asymptote, and threshold parameters.
  • The combined posterior samples from all chains.
  • Rejection rates for each of the chains (optionally included; see Table 4).
  • The input pedigree data for reference.
  • Summary statistics for the posterior distributions (optionally included; see Table 4).
  • Density plots of the posterior distributions for each parameter (optionally included; see Table 4).
  • Trace plots of the MCMC chains for each parameter (optionally included; see Table 4).
  • Autocorrelation function (ACF) plots for the posterior samples (optionally included; see Table 4).
  • Log-likelihood plots over iterations (optionally included; see Table 4).
  • Penetrance plots showing the resulting cumulative penetrance curves with credible intervals (optionally included; see Table 4).
  • Penetrance plots showing the probability density function (PDF) curves with credible intervals (optionally included; see Table 4).
Table 4: List of output options.
Option Name Value Default Value Definition
summary_stats TRUE/FALSE TRUE Standard summary statistics computed based on the posterior samples for each parameter.
rejection_rates TRUE/FALSE TRUE Rejection rates for each of the chains.
density_plots TRUE/FALSE TRUE Density plots showing the posterior distribution for each parameter.
plot_trace TRUE/FALSE TRUE Trace plots showing the MCMC sampling over iterations for each parameter.
plot_acf TRUE/FALSE TRUE Autocorrelation function (ACF) plots for the posterior samples of each parameter.
plot_loglikelihood TRUE/FALSE TRUE Plots of the log-likelihood values over MCMC iterations.
penetrance_plot TRUE/FALSE TRUE Penetrance plots showing the cumulative penetrance curves with credible intervals based on the posterior samples.
penetrance_plot_pdf TRUE/FALSE TRUE Penetrance plots showing the probability density function (PDF) curves with credible intervals based on the posterior samples.
probCI Float (0 to 1) 0.95 Probability level for credible intervals in penetrance plots.
4 Illustration
4.1 Data
To illustrate our method, we conducted a simulation study designed to mirror real-world familial colorectal cancer (CRC) patterns associated with Lynch syndrome. The simulation parameters were specifically chosen to reflect characteristics of PGVs in the MLH1 gene, one of the most common mismatch repair genes associated with Lynch Syndrome. We based our carrier prevalence on population-level MLH1 data, while the penetrance parameters (for simulating disease status) were derived from established CRC risk estimates in MLH1 carriers (Wang et al., 2020; Felton et al., 2007).
We simulated 130 probands with PGVs in the MLH1 gene, along with a total of 4,604 individuals including probands (average family size= 35.4) with 428 CRC diagnoses (203 females and 225 males). All probands and relatives had complete age information; there were no missing age values, both for age of diagnosis for affected individuals and censoring age for unaffected individuals. All probands were assumed to be carriers, and all relatives were assumed to have unknown carrier status. The dataset, code and resulting output of this illustration can be found in the vignette of the R package.
Figure 3: Exemplary code for running penetrance for simulated data for MLH1 and CRC.
4.2 Results
We run the sex-specific estimation procedure using the default settings of the penetrance package, which include non-informative priors for the Weibull parameters, a single MCMC chain, and default proposal variances (see Figure 3). The baseline risk estimates, $baseline\_data\_default$, which approximate the noncarrier risk, are CRC incidence rates from 8 registries for the US population from 1975 to 2020 as reported by SEER (Howlader et al., 2020). The user could specify alternative priors on the parameters of the Weibull distribution as discussed in the section on prior specification. Since the simulated data was generated with MLH1 carriers, we specify the carrier prevalence for MLH1 in the function call ($prev=prevMLH1$).
We generate a chain of 20,000 posterior samples and discard the first 2,000 samples as burn-in ($burn\_in=0.1$) in each plot. We do not perform any age imputation ($age\_imputation=FALSE$ and do not attempt any ascertainment correction by removing the proband ($remove\_proband=FALSE)$. Figure 4 shows the estimated absolute and cumulative penetrance curves for females and males. The shaded areas represent the 95% credible intervals derived from the posterior samples. The penetrance function returns the outputs discussed in the package section and an R object that includes the posterior samples which can be used for subsequent analysis.
Figure 4: Age-specific absolute risk (A) and cumulative risk (B) for CRC and MLH1 for females and males with default prior parameters for simulated data.
5 Discussion
Estimating disease-specific penetrance from complex family-based studies with a limited number of families poses significant challenges. However, accurate penetrance estimation remains important as the number of studies with genotyped family members continues to grow. Penetrance estimates are a critical input to risk stratification for early detection and other decision support aspects of managing inherited susceptibility of diseases including cancer.
In this work we developed an easy-to-use, well documented, open-source R package for quantitative penetrance assessment. The software offers a user-friendly approach for estimating age-specific penetrance, particularly suitable for complex family-based studies. By using Bayesian methods, the user can integrate existing information on penetrance curves through the specification of hyperparameters on the priors of the Weibull distribution parameters. We also provide uninformative default settings that allow a user to get started with minimal input while offering an extensive list of options to customize the penetrance estimation as desired.
The package includes checks on the input data and the option to automatically impute missing age data, addressing a common challenge in family-based studies. These studies often rely on family history data, typically provided by a single participant, usually the proband, who recounts the medical history of the entire family. The run-time of the algorithm increases with the age imputation. The higher the frequency of the imputation (set with the parameter $imp\_interval$), the longer the run-time. Without the age imputation the run-time scales linearly with the number of iterations and the number of families (and family size). With small numbers of families and small family sizes, the statistical reliability of the estimates decreases and credible intervals tend to be wider. Users should exercise caution when interpreting results from small samples, and it may be advisable to increase the number of iterations to better characterize the uncertainty.
Our current software implementation uses a Weibull parameterization which assumes unimodal penetrance curves, which may not fully capture more complex, multi-modal age-specific patterns. The software currently supports the joint estimation of two sex-specific penetrance functions with a single set of Weibull parameter priors, though this framework could easily be extended to accommodate different priors for each sex as well as additional stratification dimensions. Our model is limited to evaluating a single disease, such as a specific cancer, associated with a single carrier status and does not incorporate competing risks. Individuals may, for instance, develop multiple primary cancers throughout their lifetime and are subject to various other causes of death. Future work should extend this approach to richer datasets, modeling multiple cancer types simultaneously to better assess lifetime cancer risks. Furthermore our model currently assumes that homozygous carriers are not viable, which is a reasonable assumption for many rare syndromes.
In conclusion, our open-source R package represents a significant advancement in penetrance estimation for complex family-based studies. By combining robust statistical methodology with user-friendly implementation, we have created a valuable tool for genetic epidemiology research. While acknowledging current limitations, our software provides a foundation for more sophisticated penetrance estimation approaches, ultimately contributing to improved risk assessment and clinical decision support in inherited disease management. Our initial illustration in cancer genetics demonstrates the package’s utility, and we anticipate its application across a broad spectrum of hereditary conditions.
6 Acknowledgments
We gratefully acknowledge support from the National Cancer Institute at the National Institutes of Health grants R01CA242218. We thank Dr. Chris Amos for very helpful comments.
7 Data Availability
This manuscript introduces penetrance, an innovative penetrance estimation tool. The package is available on CRAN and the source code can be accessed at https://github.com/nicokubi/penetrance
Supplementary Material for: Development of the penetrance R package for Penetrance Estimation in Family-based Studies”
by N. Kubista, D. Braun, G. Parmigiani
Appendix A Prior Elicitation based on existing studies
The package provides the user with the option to automatically generate priors based on information gathered in existing penetrance studies. Since users may often not have enough knowledge to directly define the custom parameters for the priors, we provide three options to facilitate the incorporation of different types of data that a user might be able to recover from published studies:
  1. 1.
    Relative Risk Estimates: This option uses published measures of overall relative disease risk (e.g., odds ratios [OR] or relative risks [RR]) to inform the prior on the asymptote parameter, which represents lifetime penetrance. Specifically, if an overall disease risk measure for carriers is reported, you can enter it using the parameter $ratio$. The software then multiplies this ratio by the cumulative baseline risk to determine the mean of the asymptote parameter’s prior.
  2. 2.
    Granular Age-Specific Risk Data: This approach requires detailed data on the age distribution of disease diagnoses in a study. Such information may be recovered, for example, from Kaplan-Meier (K-M) curves. In the $distribution\_data\_default$ object, users must populate two rows based on the information from the study: $age$ and $at\_risk$. The $age$ row should contain four time points: the earliest age of diagnosis (row name $min$), the age at which 25% of cases occurred (row name $first\_quartile$), the age at which 50% of cases occurred (row name $median$), and the maximum age (row name $max$). The $at\_risk$ row must contain the corresponding number of carriers still under observation at each of these ages (see Figure A1 - Panel a). Based on these inputs the algorithm then computes the parameters $\alpha$ and $\beta$ for the beta distributions of the priors for the median, first quartile, and asymptote (see Table A1).
  3. 3.
    High-level Age-Specific Risk Data: This option is suitable when only the age distribution of diagnoses and the overall study size are available (but not the number of people at risk at every time point). Users populate just the $age$ column in the $distribution\_data\_default$ object with the same four time points described above (using identical row names) and provide the total number of carriers through the $sample\_size$ parameter. The package then calculates the number of at-risk carriers at each age point using predefined proportions: 90% of the total sample size at the first quartile age, 50% at the median age, and 10% at the maximum age see Figure (A1 - Panel b).
(a) Example configuration with user-provided prior information on ages and resp. number of individuals at risk at those ages.
(b) Example configuration with user-provided ages and the total sample size.
Figure A1: Structure of the input object for prior information.
Table A1: Automatic prior elicitation using user inputs in distribution data default (Setting 3).
Model Parameter Distribution Parameter 1 (Alpha) Parameter 2 (Beta)
First Quartile Scaled Beta From distribution_data_default: Normalized first quartile * number of individuals at risk at the first quartile Number of individual at risk at the first quartile - alpha
Median Scaled Beta From distribution_data_default: Normalized median * number of individuals at risk at the median Number of individual at risk at the median - alpha
Asymptote Scaled Beta From distribution_data_default: Normalized max. age * number of individuals at risk at the max. age Number of individual at risk at max. age - alpha
Threshold Uniform 0 From distribution_data_default: Min. age
Appendix B Thresholds for the Sampling of Proposals
Parameter bounds in our MCMC algorithm ensure biological plausibility while maintaining computational efficiency. These thresholds prevent the exploration of biologically impossible parameter combinations while ensuring interpretable results. The specific bounds are detailed in Table B2.
Parameter Condition
Asymptote $0\leq\text{Asymptote}\leq 1$
Threshold $0\leq\text{Threshold}\leq 100$
$Q_{50}$ (Median) $Q_{50}\geq Q_{25}$
[…]
$Q_{25}$ (First Quartile) $\text{Threshold}\leq Q_{25}\leq Q_{50}$
Table B2: Conditions for the parameters irrespective of sex.